# Two-phase flows through porous media described by a Cahn–Hilliard–Brinkman model with dynamic boundary conditions

### v1

@@ 3.1.5   Additional regularity for the phase-fields
for any $\varepsilon\in(0,1)$. Applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we infer $\big{(}\varphi_{k}(t),\psi_{k}(t)\big{)}\in H^{3}(\Omega)\times H^{3}(\Gamma)$ with
$\left\|\varphi_{k}(t)\right\|_{H^{3}(\Omega)}^{2}+\left\|\psi_{k}(t)\right\|_{H^{3}(\Gamma)}^{2}\leq C\left\|f_{k}(t)\right\|_{V}^{2}+C\left\|g_{k}(t)\right\|_{{V}_{\Gamma}}^{2}$
$\quad\leq C\left\|\mu_{k}(t)\right\|_{V}^{2}+C\left\|\theta_{k}(t)\right\|_{{V}_{\Gamma}}^{2}+C\varepsilon^{-2}+C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{2}+C\varepsilon^{2}\|\psi_{k}(t)\|_{H^{3}(\Gamma)}^{2}.$
After choosing $\varepsilon\in(0,1)$ sufficiently small, we absorb the terms $C\varepsilon^{2}\|\varphi_{k}(t)\|_{H^{3}(\Omega)}^{2}$ and $C\varepsilon^{2}\|\psi_{k}(t)\|_{H^{3}(\Gamma)}^{2}$ by the left-hand side. Integrating the resulting estimate with respect to $t$ from 00 to $T$, we use (3.1.3) to eventually obtain the uniform estimate
$\left\|\varphi_{k}\right\|_{L^{2}(0,T;H^{3}(\Omega))}+\left\|\psi_{k}\right\|_{L^{2}(0,T;H^{3}(\Gamma))}\leq C.$ (3.46)
Therefore, after another subsequence extraction, we have, as $k\to\infty$,
$\varphi_{k}$ $\to\varphi\quad\text{weakly in $L^{2}(0,T;H^{3}(\Omega))$},$
$\psi_{k}$ $\to\psi\quad\text{weakly in $L^{2}(0,T;H^{3}(\Gamma))$}.$
Hence, the regularity assertion (2.15b) is verified, and the regularity (2.16) follows directly from Proposition A.1.
@@
This means that all claims are established and thus, the proof is complete. ∎
@@ 3.2   The limit $K\to 0$ and existence of a weak solution in the case $K=0$
for all $t\in[0,T]$ and all $n\in{\mathbb{N}}$. Hence, proceeding analogously as in the proof of Theorem 2.4 (Subsection 3.1.2, First and Second estimates), we derive the uniform bound
@@
$\left|\int_{\Omega}\mu^{K_{n}}\eta\right|\leq\left\|\varphi^{K_{n}}\right\|_{V}\left\|\eta\right\|_{V}+\left\|F^{\prime}(\varphi^{K_{n}})\right\|_{L^{\frac{6}{5}}(\Omega)}\left\|\eta\right\|_{L^{6}(\Omega)}\leq C\left\|\eta\right\|_{V}$ (3.50)
entailing that $\mu^{K_{n}}$ is bounded in $L^{\infty}(0,T;V^{*})$. Then, by means of the generalized Poincaré inequality (see [8, Section 8.16] or [33, Lemma A.1]), we infer that
$\left\|\mu^{K_{n}}\right\|_{H}\leq C\big{(}1+\left\|\nabla\mu^{K_{n}}\right\|_{\bf{H}}\big{)}.$
For more details see, e.g., [33, Proof of Lemma 4.5]. Hence, in combination with (3.50), we conclude
@@
Employing Poincaré’s inequality, we thus infer
@@
In combination with (3.48), this proves that $\varphi^{0}|_{\Gamma}=\psi^{0}$ a.e. on $\Sigma$ due to uniqueness of the limit. Proceeding further as in Subsection 3.1.4, we eventually show that the quintuplet $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman system (1) in the sense of Definition 2.3.
It remains to verify the additional regularity assertions. Without loss of generality, we merely consider the case $d=3$. The case $d=2$ can be handled analogously but is even easier as the Sobolev embeddings in two dimensions are better. We infer from (2.13d) written for the solution $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ and $K=0$ that there exists a null set $\mathcal{N}\subset[0,T]$ such that
@@
$-\Delta\varphi^{0}(t)$ $=f(t)$ in $\quad\text{in $\Omega$},$ (3.57a)
$-\Delta_{\Gamma}\psi^{0}(t)+\partial_{\bf n}\varphi^{0}(t)$ $=g(t)$ on $\quad\text{on $\Gamma$},$ (3.57b)
$\varphi^{0}(t)|_{\Gamma}$ $=\psi^{0}(t)$ on $\quad\text{on $\Gamma$},$ (3.57c)
@@
$\left\|F^{\prime}\big{(}\varphi^{0}(t)\big{)}\right\|_{H}$ $\leq C+C\left\|\varphi^{0}\right\|_{L^{6}(\Omega)}^{3}\leq C,$ (3.58)
$\left\|G^{\prime}\big{(}\psi^{0}(t)\big{)}\right\|_{{H}_{\Gamma}}$ $\leq C+C\left\|\psi^{0}\right\|_{L^{2(q-1)}(\Gamma)}^{q-1}\leq C.$ (3.59)
Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we find that $\big{(}\varphi^{0}(t),\psi^{0}(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
@@
$\quad\leq C+C\left\|\mu^{0}(t)\right\|_{H}^{2}+C\left\|\theta^{0}(t)\right\|_{{H}_{\Gamma}}^{2}.$
Since $\mu^{0}\in L^{2}(0,T;H)$ and $\theta^{0}\in L^{2}(0,T;{H}_{\Gamma})$, this proves (2.18a).
$\left\|F^{\prime\prime}\big{(}\varphi^{0}(t)\big{)}\nabla\varphi^{0}(t)\right\|_{\bf H}$ $\leq C\left\|\nabla\varphi^{0}(t)\right\|_{{\bf H}}+C\left\||\varphi^{0}(t)|^{2}\,\nabla\varphi^{0}(t)\right\|_{{\bf H}}$
$\leq C+C\left\|\varphi^{0}(t)\right\|_{L^{6}(\Omega)}^{2}\,\left\|\nabla\varphi^{0}(t)\right\|_{L^{6}(\Omega)}$
$\leq C+C\left\|\varphi^{0}(t)\right\|_{H^{2}(\Omega)}$
and
$\left\|G^{\prime\prime}\big{(}\psi^{0}(t)\big{)}\nabla_{\Gamma}\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}$ $\leq C\left\|\nabla\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}+C\left\||\psi^{0}(t)|^{q-2}\,\nabla\psi^{0}(t)\right\|_{{\bf H}_{\Gamma}}$
$\leq C+C\left\|\psi^{0}(t)\right\|_{L^{{2(q-2)}}}^{q-2}\left\|\nabla\psi^{0}(t)\right\|_{L^{4}(\Gamma)}$
$\leq C+C\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}.$
In combination with (3.58) and (3.59), these estimates directly imply
$\left\|F^{\prime}\big{(}\varphi^{0}(t)\big{)}\right\|_{V}$ $\leq C+C\left\|\varphi^{0}\right\|_{H^{2}(\Omega)},$
$\left\|G^{\prime}\big{(}\psi^{0}(t)\big{)}\right\|_{{V}_{\Gamma}}$ $\leq C+C\left\|\psi^{0}(t)\right\|_{H^{2}(\Gamma)}.$
Now, applying regularity theory for elliptic problems with bulk-surface coupling (see [48, Theorem 3.3]), we infer
@@
Recalling $\mu^{0}\in L^{2}(0,T;V)$, $\theta^{0}\in L^{2}(0,T;{V}_{\Gamma})$ and that (2.18a) with $\ell=2$ is already verified, this proves (2.18b) in the case $\ell=3$. By means of Proposition A.1(b), we directly infer $(\varphi^{0},\psi^{0})\in C^{0}([0,T];{\cal V}_{0})$. Moreover, via interpolation between $L^{\infty}(0,T;{\cal V}_{0})$ and $L^{2}(0,T;H^{3}(\Omega)\times H^{3}(\Gamma))$ (cf. Lemma 2.2), we further get
@@ 3.3   Uniqueness of the weak solution for regular potentials
We now intend to control the terms $I_{i}$, $i=1,...,10$, by means of Hölder’s inequality, Young’s inequality, the Lipschitz continuity of $\lambda$ and $\gamma$, and integration by parts along with Sobolev’s embeddings and the trace theorem. For a positive $\delta$ yet to be chosen, we derive the following estimates:
@@
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{s}(\Omega)}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{2p-8}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ for $\text{for $d=2$},$
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{(p-2)}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ for $\text{for $d=3$},$
$\left\|\varphi_{2}\right\|_{W^{1,4}(\Omega)}^{2}\leq C\left\|\varphi_{2}\right\|_{H^{1}(\Omega)}^{\frac{1}{2}}\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}\leq C\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}$ for $\text{for $d=2,3$},$
@@
In the case $d=3$, we use interpolation between Sobolev spaces (Lemma 2.2) to derive the embedding
@@
$\varphi_{i}\in L^{16}(0,T;L^{12}(\Omega))\cap L^{\frac{48}{5}}(0,T;L^{36}(\Omega)),\quad i=1,2,$
by choosing $\rho=16$ and $\rho=\frac{48}{5}$ in (3.67), respectively.
@@
In view of (2.21), we assume, without loss of generality, that $q\geq 5$. Recalling that the boundary $\Gamma$ is a $(d-1)$-dimensional submanifold of ${\mathbb{R}}^{d}$ with $d\in\{2,3\}$, we have $H^{s}(\Gamma){\hookrightarrow}L^{\infty}(\Gamma)$ for every $s>1$. Hence, via interpolation between Sobolev spaces (see Lemma 2.2) we obtain the estimate
@@
$\delta\max\Big{\{}4,C,C\left\|\varphi_{2}\right\|_{L^{\infty}(0,T;V)}^{2},C\left\|\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}^{2}\Big{\}}\leq\frac{1}{4}\min\Big{\{}{{M}},{{M_{\Gamma}}},{C_{P}(\nu,\gamma_{1})}\Big{\}},$
where $C_{P}(\nu,\gamma_{1})$ is a Poincaré constant such that $2\nu\left\|\mathbf{D}{\boldsymbol{v}}\right\|^{2}_{\mathbb{H}}+\gamma_{1}\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf H}_{\Gamma}}\geq C_{P}(\nu,\gamma_{1})\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf{V}}}$. Thus, we integrate over time and employ Gronwall’s lemma to deduce that

### v2

@@ 3.1.5 Additional regularity for the phase-fields
Hence, applying regularity theory for elliptic problems with bulk-surface coupling (see [61, Theorem 3.3]), we find that $\big{(}\varphi(t),\psi(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
$\left\|\varphi(t)\right\|_{H^{2}(\Omega)}^{2}+\left\|\psi(t)\right\|_{H^{2}(\Gamma)}^{2}\leq C\left\|f(t)\right\|_{H}^{2}+C\left\|g(t)\right\|_{{H}_{\Gamma}}^{2}$
$\quad\leq C+C\left\|\mu(t)\right\|_{H}^{2}+C\left\|\theta(t)\right\|_{{H}_{\Gamma}}^{2}.$
Since $\mu\in L^{4}(0,T;H)$ and $\theta\in L^{4}(0,T;{H}_{\Gamma})$, this proves (2.19a).
$\left\|F^{\prime\prime}\big{(}\varphi(t)\big{)}\nabla\varphi(t)\right\|_{\bf{H}}$ $\leq C\left\|\nabla\varphi(t)\right\|_{{\bf{H}}}+C\left\||\varphi(t)|^{2}\,\nabla\varphi(t)\right\|_{{\bf{H}}}$
$\leq C+C\left\|\varphi(t)\right\|_{L^{6}(\Omega)}^{2}\,\left\|\nabla\varphi(t)\right\|_{{{\bf{L}}^{6}(\Omega)}}$
$\leq C+C\left\|\varphi(t)\right\|_{H^{2}(\Omega)}$
and
$\left\|G^{\prime\prime}\big{(}\psi(t)\big{)}\nabla_{\Gamma}\psi(t)\right\|_{{\bf{H}}_{\Gamma}}$ $\leq C\left\|\nabla\psi(t)\right\|_{{\bf{H}}_{\Gamma}}+C\left\||\psi(t)|^{q-2}\,\nabla\psi(t)\right\|_{{\bf{H}}_{\Gamma}}$
$\leq C+C\left\|\psi(t)\right\|_{L^{{2(q-2)}}}^{q-2}\left\|\nabla\psi(t)\right\|_{{{\bf{L}}^{4}(\Gamma)}}$
$\leq C+C\left\|\psi(t)\right\|_{H^{2}(\Gamma)}.$
In combination with (3.44) and (3.45), these estimates directly imply
$\left\|F^{\prime}\big{(}\varphi(t)\big{)}\right\|_{V}\leq C+C\left\|\varphi(t)\right\|_{H^{2}(\Omega)},\quad\left\|G^{\prime}\big{(}\psi(t)\big{)}\right\|_{{V}_{\Gamma}}\leq C+C\left\|\psi(t)\right\|_{H^{2}(\Gamma)}.$
Now, applying regularity theory for elliptic problems with bulk-surface coupling (see [61, Theorem 3.3]), we infer
$\left\|\varphi(t)\right\|_{H^{3}(\Omega)}^{2}+\left\|\psi(t)\right\|_{H^{3}(\Gamma)}^{2}\leq C\left\|f(t)\right\|_{V}^{2}+C\left\|g(t)\right\|_{{V}_{\Gamma}}^{2}$
$\quad\leq C+C\left\|\mu(t)\right\|_{V}^{2}+C\left\|\theta(t)\right\|_{{V}_{\Gamma}}^{2}+C\left\|\varphi(t)\right\|_{H^{2}(\Omega)}^{2}+C\left\|\psi(t)\right\|_{H^{2}(\Gamma)}^{2}.$
Recalling $\mu\in L^{2}(0,T;V)$, $\theta\in L^{2}(0,T;{V}_{\Gamma})$ and that (2.19a) with $\ell=2$ is already verified, this proves (2.19b). By means of Proposition A.1(b), we directly infer $(\varphi,\psi)\in C^{0}([0,T];{\cal V}_{0})$, which proves (2.20).
@@
This means that all assertions are verified and thus, the proof is complete.
@@ 3.2  The limit $K\to 0$ and existence of a weak solution in the case $K=0$
for all $t\in[0,T]$ and all $n\in{\mathbb{N}}$. Hence, proceeding similarly as in the proof of Theorem 2.6 (Subsection 3.1.2, First and Second estimates), we derive the uniform bound
@@
$\left|\int_{\Omega}\mu^{K_{n}}\eta\right|\leq\left\|\varphi^{K_{n}}\right\|_{V}\left\|\eta\right\|_{V}+\left\|F^{\prime}(\varphi^{K_{n}})\right\|_{L^{6/5}(\Omega)}\left\|\eta\right\|_{L^{6}(\Omega)}\leq C\left\|\eta\right\|_{V}.$ (3.49)
Fixing a function $\eta\in C^{\infty}_{c}(\Omega)$ with ${\langle\eta\rangle}_{\Omega}\neq 0$, we deduce
$\left\|\mu^{K_{n}}\right\|_{H}\leq C\big{(}1+\left\|\nabla\mu^{K_{n}}\right\|_{\bf{H}}\big{)}$
by means of a generalized Poincaré inequality (see, e.g., [36, Lemma B.63]). Hence, in combination with (3.49), we conclude
@@
Employing Poincaré’s inequality on $\Gamma$ (see Lemma 2.2), we thus infer
@@
In combination with (3.47), this proves that $\varphi^{0}|_{\Gamma}=\psi^{0}$ a.e. on $\Sigma$ due to uniqueness of the limit. Proceeding further as in Subsection 3.1.4, we eventually show that the quintuplet $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ is a weak solution of the Cahn–Hilliard–Brinkman system (1.1) in the sense of Definition 2.4.
It remains to verify the additional regularity assertions. Therefore, we proceed similarly as in Subsection 3.1.5. Without loss of generality, we merely consider the case $d=3$. The case $d=2$ can be handled analogously but is even easier as the Sobolev embeddings in two dimensions are better. We infer from (2.16d) written for the solution $({\boldsymbol{v}}^{0},\varphi^{0},\mu^{0},\psi^{0},\theta^{0})$ and $K=0$ that there exists a null set $\mathcal{N}\subset[0,T]$ such that
@@
$-\Delta\varphi^{0}(t)$ $=f(t)$ $\quad\text{in $\Omega$},$ (3.56a)
$-\Delta_{\Gamma}\psi^{0}(t)+\partial_{\bf n}\varphi^{0}(t)$ $=g(t)$ $\quad\text{on $\Gamma$},$ (3.56b)
$\varphi^{0}(t)|_{\Gamma}$ $=\psi^{0}(t)$ $\quad\text{on $\Gamma$},$ (3.56c)
@@
Let us first consider the case $\ell=2$. Proceeding exactly as in Subsection 3.1.5, we deduce that $\big{(}\varphi^{0}(t),\psi^{0}(t)\big{)}\in H^{2}(\Omega)\times H^{2}(\Gamma)$ with
@@
$\quad\leq C+C\left\|\mu^{0}(t)\right\|_{H}^{2}+C\left\|\theta^{0}(t)\right\|_{{H}_{\Gamma}}^{2}$
thanks to regularity theory for elliptic problems with bulk-surface coupling (see [61, Theorem 3.3]). Since $\mu^{0}\in L^{2}(0,T;H)$ and $\theta^{0}\in L^{2}(0,T;{H}_{\Gamma})$, this proves (2.22a).
We now consider the case $\ell=3$. Proceeding analogously as in Subsection 3.1.5, we infer
@@
by means of regularity theory for elliptic problems with bulk-surface coupling (see [61, Theorem 3.3]). Recalling $\mu^{0}\in L^{2}(0,T;V)$, $\theta^{0}\in L^{2}(0,T;{V}_{\Gamma})$ and that (2.22a) with $\ell=2$ is already verified, this proves (2.22b). With the help of Proposition A.1(b), we directly infer $(\varphi^{0},\psi^{0})\in C^{0}([0,T];{\cal V}_{0})$. Moreover, via interpolation between $L^{\infty}(0,T;{\cal V}_{0})$ and $L^{2}(0,T;H^{3}(\Omega)\times H^{3}(\Gamma))$ (cf. Lemma 2.3), we further get
@@ 3.3  Uniqueness of the weak solution for regular potentials
We now intend to control the terms $I_{i}$, i=1,,10i=1,...,10, by means of Hölder’s inequality, Young’s inequality, the Lipschitz continuity of $\lambda$ and $\gamma$, and integration by parts along with Sobolev’s embeddings and the trace theorem. For a positive $\delta$ yet to be chosen, we derive the following estimates:
@@
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{s}(\Omega)}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{2p-8}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ $\text{for $d=2$},$
$\left\|\varphi_{1}\right\|_{{L^{\infty}(\Omega)}}^{2(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{1}(\Omega)}^{(p-2)}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{(p-2)}\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{4}$ $\text{for $d=3$},$
$\left\|\varphi_{2}\right\|_{W^{1,4}(\Omega)}^{2}\leq C\left\|\varphi_{2}\right\|_{H^{1}(\Omega)}^{\frac{1}{2}}\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}\leq C\left\|\varphi_{2}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}$ $\text{for $d=2,3$},$
@@
In the case $d=3$, we use interpolation between Sobolev spaces (Lemma 2.3) along with the continuous embedding $H^{(4+\rho)/\rho}(\Omega){\hookrightarrow}L^{6\rho/(\rho-8)}(\Omega)$ to derive the estimate
$\int_{0}^{T}\|u(t)\|_{L^{6\rho/(\rho-8)}(\Omega)}^{\rho}$ $\leq C\int_{0}^{T}\|u(t)\|_{H^{(4+\rho)/\rho}(\Omega)}^{\rho}\leq C\int_{0}^{T}\|u(t)\|_{H^{3}(\Omega)}^{2}\|u(t)\|_{H^{1}(\Omega)}^{\rho-2}$
$\leq C\|u\|_{L^{2}(0,T;H^{3}(\Omega))}^{2}\|u\|_{L^{\infty}(0,T;H^{1}(\Omega))}^{\rho-2}$
for any $\rho>8$ and any function $u\in L^{\infty}(0,T;H^{1}(\Omega))\cap L^{2}(0,T;H^{3}(\Omega))$. This proves the continuous embedding
@@
$\varphi_{i}\in L^{16}(0,T;L^{12}(\Omega)),\quad i=1,2,$
by choosing $\rho=16$ in (3.64).
@@
In view of (2.25), we assume, without loss of generality, that $q\geq 5$. Recalling that the boundary $\Gamma$ is a (d1)(d-1)-dimensional submanifold of ${\mathbb{R}}^{d}$ with $d\in\{2,3\}$, we have $H^{s}(\Gamma){\hookrightarrow}L^{\infty}(\Gamma)$ for every $s>1$. Hence, via interpolation between Sobolev spaces (see Lemma 2.3) we obtain the estimate
@@
$\delta\max\Big{\{}4,C,C\left\|\varphi_{2}\right\|_{L^{\infty}(0,T;V)}^{2},C\left\|\psi_{2}\right\|_{L^{\infty}(0,T;{V}_{\Gamma})}^{2}\Big{\}}\leq\frac{1}{4}\min\Big{\{}{{M}},{{M_{\Gamma}}},{C_{\star}(\nu,\gamma_{1})}\Big{\}}.$
Here, the constant $C_{\star}(\nu,\gamma_{1})$ results from Korn’s inequality (see Lemma 2.1) and is chosen such that $2\nu\left\|\mathbf{D}{\boldsymbol{v}}\right\|^{2}_{\mathbb{H}}+\gamma_{1}\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf{H}}_{\Gamma}}\geq C_{\star}(\nu,\gamma_{1})\left\|{\boldsymbol{v}}\right\|^{2}_{{\bf{V}}}$. Thus, we integrate over time and employ Gronwall’s lemma to deduce that